# Co-Designing Multimodal Systems for Accessible Remote Dance Instruction

### v1

@@ Abstract.
Videos make exercise instruction widely available, but they rely on visual demonstrations that blind and low vision (BLV) learners cannot see. While audio descriptions (AD) can make videos accessible, describing movements remains challenging as the AD must convey what to do (mechanics, location, orientation) and how to do it (speed, fluidity, timing). Prior work thus used multimodal instruction to support BLV learners with individual simple movements. However, it is unclear how these approaches scale to dance instruction with unique, complex movements and precise timing constraints. To inform accessible remote dance instruction systems, we conducted three co-design workshops (N=28) with BLV dancers, instructors, and experts in sound, haptics, and AD. Participants designed 8 systems revealing common themes: staged learning to dissect routines, crafting vocabularies for movements, and selectively using modalities—narration for movement structure, sound for expression, and haptics for spatial cues. We conclude with design recommendations to make learning dance accessible.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Accessibility
@@ 1. Introduction
Millions of people now participate in exercises such as dance, yoga, and fitness using online videos. As exercise videos allow people to learn on their own time and in their own environment, their popularity expanded with the COVID-19 pandemic (Nyberg et al., 2024; Sui et al., 2022; Delabary et al., 2022; Schmid and and, 2022) and the corresponding rise in the use of video platforms (Mitkina et al., 2023; Nuhn et al., 2025; Warburton, 2024; Hong et al., 2020). Such videos also have the potential to be particularly useful for people with disabilities as remote learning reduces travel barriers associated with in-person physical instruction (Jaarsma et al., 2014; lee, 2014), and increased physical activity can improve life satisfaction (Łabudzki and Tasiemski, 2013). Yet, this revolution in accessible fitness has systematically excluded blind and low vision (BLV) learners, as instructors primarily rely on visual demonstrations (Bläsing et al., 2018) that BLV learners cannot see.
@@
aPrior work thus explored how to provide accessible feedback to BLV learners with multi-modal approaches — from verbal feedback to correct yoga pose positioning (Rector et al., 2013) to haptic and audio feedback to improve performance in tennis (Morelli et al., 2010b) and bowling (Morelli et al., 2010a) exergames. Such tools provide feedback loops that can improve performance and confidence (Nikooyan and Ahmed, 2015; Todorov et al., 1997), and alternative modalities (sound, haptic, tactile) that can support a more complete understanding of movements when combined with verbal narrations (Bläsing and Zimmermann, 2021; Seham and and, 2015; and, 2023). Dance movements in particular can be challenging to convey and correct due to the diverse, complex positioning of the body, such that prior work explored how to make movement instruction more accessible for in-person Contemporary dance classes with multiple modalities (De Silva et al., 2025). While this rich body of prior work demonstrates that communicating and correcting individual movements to BLV learners is possible, it is not clear how to scale such knowledge to long, multi-move routines of online videos without the support of co-located human instruction and feedback.
@@
In this work, we use a participatory design approach to investigate how to make remote dance learning accessible to BLV learners by including BLV learners and dance instructors as equal designers rather than mere consumers of technology. We select dance as our domain of interest due to the rich complexity in movement mechanics, timing, and multi-movement sequences that makes dance particularly challenging to learn. To co-design systems for remote dance instruction, we conduct co-design workshops with participants that have a wide range of relevant expertise to draw from complementary perspectives that are relevant to learning, instructing, and communicating around dance: BLV dance learners, dance instructors, and technical experts (in AD, sound, and haptics), along with participants with multiple types of expertise (e.g., BLV dance instructors). Our workshops aimed to answer the research question: How should we design systems to support remote dance instruction for BLV learners?
@@
Over the course of three co-design workshops with eight design groups, 28 total participants generated considerations for accessible remote dance instructions and created eight unique system designs. Our analysis revealed several common themes across system architectures: using staged learning to disentangle placement and timing concerns, breaking routines into small movements, crafting multimodal vocabularies for movements (e.g., unique names, or sound cues), and selectively using multiple modalities across learning stages — e.g., narration for movement mechanics, sound for conveying expression, and haptics for spatial cues. We conclude with design guidelines for accessible learning systems to inform future system design and reserach.
@@ 2.2. Accessible Movement Instruction
Exergames, or video games that integrate physical exercises into engaging experiences, often combine modalities that can increase engagement without vision. For instance, audio+haptic tennis (Morelli et al., 2010b) and bowling (Morelli et al., 2010a) have been shown to improve guidance and enjoyment. Prior works have also explored sensor-based systems for home exercise (Malik et al., 2021; Rector et al., 2013), yet they focus on guiding stance and alignment by evaluating discrete poses, rather than the continuous transitions or expression central to dance. Feedback is typically stepwise (e.g., verbal prompts, success tones) and correctness is judged at end-states (e.g., hit/hold/balance), not on how motions unfold or whether movement qualities align with intent. Dance learning from video, by contrast, demands precise timing, nuanced transitions between shapes, and attention to expressive dynamics. Our work addresses this gap by exploring accessible dance education in  remote video settings, which present unique opportunities and considerations for multimodal solutions.
@@ 2.3. Multi-Modal Feedback for Accessibility
Figure 1. Our co-design workshop featured 4 major phases. The participants were informed of the workshop goals and their task (Introduction), then they split into smaller groups to teach the dance clips to one another (Teaching), design a system to effectively teach the clip using what they learned from teaching one another (Designing) and then shared their systems with all the other groups (Presentation).
@@ 3. Method
    What are the current needs and design considerations of BLV learners seeking to learn dance on their own?
@@
    How can we design accessible systems for dance learning with verbal narration, non-verbal audio, and/or haptics?
To incorporate the deep expertise and lived experience of relevant experts, we selected to use co-design workshops — a participatory design approach that invites community members (e.g., BLV learners, dance instructors) to be creators and originators of the technology that they use rather than only end consumers (Magnusson et al., 2018). Similar to prior co-design workshops (Magnusson et al., 2018; Aflatoony and Lee, 2020; Valencia et al., 2021), we aimed to incorporate perspectives of not only potential end consumers of the technology (BLV dance learners) but also relevant experts in dance instruction (BLV and sighted dance instructors), and experts in multimodal technologies (audio description, sound design, haptics) to build on their complementary expertise.
@@ 3.1. Workshop Design
Our in-person co-design workshops featured five stages (Figure 2): (1) a pre-workshop survey to surface relevant individual dance backgrounds and lived experiences, (2) a large group introduction to share survey findings and discuss modalities for dance instruction, (3) small groups teaching example dances using non-visual techniques, then (4) designing systems for non-visual dance instruction, and (5) presenting their systems to back to the large group for discussion. The four in-person workshop phases are shown in Figure 1. Each in-person workshop was three hours long, and we compensated participants \$90 for their participation. This study was approved by our institution’s Institutional Review Board (IRB).
@@
We began recruitment by reaching out to dance professors and students with teaching experience at our institution. Next, we connected with local BLV communities, and freelance sound producers and dance instructors. Additionally, we were able to partner with a nearby arts and accessibility organization and bring in two of their directors. Our three workshops included a total of 28 participants (16 female, 11 male, 1 non-binary), including 13 BLV participants (B1-B8b), 11 dance teachers (D1-D8b), two sound designers (S2, S6), a haptics engineer (H1), and an audio describer (A8) (Table 1). Many participants had combined expertise: 11 of 13 BLV participants had prior dance experience, B3b, D4a and H1 also had sound design expertise, and B2a, B3a, and B7b also had dance instruction expertise.
@@
Table 1. Participant demographic and background information, including gender, age, current vision, age of onset if blind, occupation, skill set (D = Dance, DI = Dance instruction, S = Sound design, H = Haptics engineering, A = Audio description), years of experience of the skill set, and dance styles if prior dance experience.
@@ 3.1.2. Workshop Stages
We designed the workshop stages to maximize contributions by all workshop members, including using accessible workshop best practices (e.g., all materials accessible, ask for accommodations in advance, provide multiple opportunities to update access requests) and using a mix of individual, small group, and large group activities. We structured large group activities to encourage participation from every workshop member (e.g., think-pair-share, explicit turn order) and we selected small groups to represent diverse expertise to foster complementary contributions. These stages are shown in Figure 2.
Figure 2. Our workshop followed a five-stage process: 1) pre-workshop survey, 2) co-design workshop introduction, 3) teaching, 4) designing, and 5) presenting
@@
From the survey results and our review of prior literature (Bläsing and Zimmermann, 2021; De Silva et al., 2023, 2025) we distilled three main themes for accessible remote dance learning to address during the co-design workshop: (1) clarity, as visual demonstrations are often not fully supplemented by other sensory forms, (2) feedback, as remote learners are currently unable to have their movements corrected by visual comparison or touch, and (3) temporal and emotional qualities, as existing verbal descriptions insufficiently convey fluidity, weight, timing, emotions, and textures in movements that are important for learning dance.
@@
We first asked all participants and research staff to take turns introducing themselves to the group and answering the question “How would you describe what makes dance meaningful to you?” to consider aspects of dance they wanted to preserve in their later design exercise. We then shared survey themes around the benefits and challenges for accessible remote movement education.
@@
Teaching & Designing: 85 minutes. Following the introductory presentation, we split participants into groups of 3-5 people with at least one BLV participant and one dance instructor per group (Table 2). We assigned each group a dance tutorial clip around 2.5 minutes long as a reference point for creating concrete designs of dance tutorial systems. We typically formed groups such that multiple participants shared experience in a common dance style, and we selected the tutorial clip to match the dance style experience of the group. We also ensured that selected clips were narrated and suitable for beginners. Partcipants remained with only their groups for the teaching and designing stages.
@@ 3 G8 B8a, B8b, D8a, D8b, A8 Cha-cha
Within their small groups, participants then discussed and designed a non-visual tutorial system for remote dance education. We also asked participants to create a concrete prototype or a walkthrough of a non-visual tutorial using their given clip as an example. We provided participants with several prototyping materials to facilitate the designing stage (Figure 3): recording devices with raised buttons for recording verbal narration and non-verbal sound cues; a custom accessible soundboard application that allowed preview and adjustment of fluid/non-fluid music, natural sounds, and synthetic sounds (as informed by movement sonification literature (Frid et al., 2018; Landry and Jeon, 2020)); 3D movable human figures for demonstrating poses; tactile stickers for prototyping haptic device placement on 3D figures or participants for pose instruction. Figure 4 shows participants using these prototyping tools to brainstorm and design their systems.
@@
Presenting: 30 minutes. Following group design sessions, all groups reconvened and were each asked to explain their non-visual tutorial system workflow in-depth, shown in Figure 5. We iterated on the workshop design following Workshop 1 and asked participants to additionally physically re-enact their system. Each group designated one or more participants to act as the system, providing instruction, haptic and tactile feedback through (with permission) physical touch, and sound cues through the soundboard and onomatopoeia. Groups designated one or more participants to act as the learners from the system. Participants narrated the system interactions as they re-enacted a tutorial for their dance clip.
@@ 4. Results
In this section, we share participants’ strategies and challenges of current dance instruction, their preferred approaches for using different modalities in remote instruction, and their concepts for future multimodal systems for end-to-end remote dance instruction.
@@ 4.1. Current Strategies and Challenges
All groups watched the assigned example dance clips during the teaching phase to learn the sequence of movements, and all groups identified major challenges with directly following the videos alone: lack of context, lack of detail, and vague terminology. For example, G8’s clip directly began demonstrating the cha-cha routine without giving context such as how many steps and movements there were. All clips did not describe orientations of body parts with enough detail to recreate the movements. G8’s clip lacked descriptiveness in body positioning and clarity with directions, such as stating to “move to the side” without mentioning left or right. B8a expressed: “It’s not clear what step she is doing or if she is stepping at all. She could be stepping, or just shifting weight.”

### v3

# Co-Designing Multimodal Systems for Accessible Asynchronous Dance Instruction

@@ Abstract.
Videos make exercise instruction widely available, but they rely on visual demonstrations that blind and low vision (BLV) learners cannot see. While audio descriptions (AD) can make videos accessible, describing movements remains challenging as the AD must convey what to do (mechanics, location, orientation) and how to do it (speed, fluidity, timing). Prior work thus used multimodal instruction to support BLV learners with individual simple movements. However, it is unclear how these approaches scale to dance instruction with unique, complex movements and precise timing constraints. To inform accessible asynchronous dance instruction systems, we conducted three co-design workshops (N=28) with BLV dancers, instructors, and experts in sound, haptics, and AD. Participants designed 8 systems revealing common themes: staged learning to dissect routines, crafting vocabularies for movements, and selectively using modalities—narration for movement structure, sound for expression, and haptics for spatial cues. We conclude with design implications to make learning dance accessible.
@@
$c$s: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Accessibility
@@ 1. Introduction
Millions of people now participate in exercises such as dance, yoga, and fitness using online videos. Exercise videos allow people to learn on their own time and in their own environment. Thus, their popularity expanded with the COVID-19 pandemic (Sui et al., 2022; Delabary et al., 2022; Suzuki and Ochiai, 2022; Chen et al., 2020) and the corresponding rise in the use of video platforms (Mitkina et al., 2023; Nuhn et al., 2025; Warburton, 2024; Hong et al., 2020). Such videos also have the potential to be particularly useful for people with disabilities as remote learning reduces travel barriers associated with in-person physical instruction (Jaarsma et al., 2014; lee, 2014), and increased physical activity can improve life satisfaction (Łabudzki and Tasiemski, 2013). Yet, this revolution in accessible fitness has systematically excluded blind and low vision (BLV) learners, as exercise videos rely on visual demonstrations (De Silva et al., 2023) that BLV learners cannot see. Compared to synchronous settings, asynchronous learning from online videos poses challenges to BLV learners given the lack of feedback and instructor adaptation (Jiang et al., 2025).
@@
Prior work thus explored how to provide accessible feedback to BLV learners with multi-modal approaches — from verbal feedback to correct yoga pose positioning (Rector et al., 2013) to haptic and audio feedback to improve performance in tennis (Morelli et al., 2010b) and bowling (Morelli et al., 2010a) exergames. Such tools provide feedback loops that can improve performance and confidence (Nikooyan and Ahmed, 2015; Todorov et al., 1997), and alternative modalities (sound, haptic, tactile) that can support a more complete understanding of movements when combined with verbal narrations (Bläsing and Zimmermann, 2021; Seham and and, 2015; and, 2023). Dance movements in particular can be challenging to convey and correct due to the diverse, complex positioning of the body, such that prior work explored how to make movement instruction more accessible for in-person Contemporary dance classes with multiple modalities (De Silva et al., 2025). While this rich body of prior work demonstrates that communicating and correcting individual movements to BLV learners is possible, it is not clear how to scale such knowledge to long, multi-move routines of online videos without the support of co-located human instruction and feedback. Our work explores creating systems to translate static instructions from asynchronous videos into accessible instruction systems for BLV learners.
@@
We use a participatory design approach to investigate how to make asychronous dance instruction accessible to BLV learners by including BLV learners and dance instructors as equal designers rather than mere consumers of technology. As we aim to maintain the advantages of asynchronous learning from dance tutorial videos available online, we define the solution that our work explores as asychronous dance instruction systems for BLV learners. We select dance as our domain of interest due to the rich complexity in movement mechanics, timing, and multi-movement sequences that makes dance particularly challenging to learn. To co-design asynchronous dance instruction systems, we conduct co-design workshops with participants that have a wide range of relevant expertise in learning, instructing, and communicating around dance: BLV dance learners, dance instructors, technical experts (in AD, sound, and haptics), and participants with multiple types of expertise (e.g., BLV dance instructors). Our workshops aimed to answer the research question: How should we design systems to support asynchronous dance instruction for BLV learners?
@@
Over the course of three in-person co-design workshops with eight design groups, 28 total participants generated considerations for accessible asynchronous dance instruction systems and created eight unique system designs. Our analysis revealed several common themes across these system designs: using staged learning to disentangle placement and timing concerns, breaking routines into small movements, crafting multimodal vocabularies for movements (e.g., unique names, or sound cues), and selectively using multiple modalities across learning stages — e.g., narration for movement mechanics, sound for conveying expression, and haptics for spatial cues. We conclude with system architectures and design implications surfaced through co-design for accessible asynchronous dance learning systems to inform future system design and research.
@@ 2.2. Accessible Movement Instruction
Exergames, or video games that integrate physical exercises into engaging experiences, often combine modalities that can increase engagement without vision. For instance, audio+haptic tennis (Morelli et al., 2010b) and bowling (Morelli et al., 2010a) have been shown to improve guidance and enjoyment. Prior works have also explored sensor-based systems for home exercise (Malik et al., 2021; Rector et al., 2013), yet they focus on guiding stance and alignment by evaluating discrete poses, rather than the continuous transitions or expression central to dance. Feedback is typically stepwise (e.g., verbal prompts, success tones) and correctness is judged at end-states (e.g., hit/hold/balance), not on how motions unfold or whether movement qualities align with intent. Dance learning from video, by contrast, demands precise timing, nuanced transitions between shapes, and attention to expressive dynamics. Our work addresses this gap by exploring accessible dance education in  asynchronous video settings, which present unique opportunities and considerations for multimodal solutions.
@@ 2.3. Multi-Modal Feedback for Accessibility
Figure 1. Our co-design workshops featured 4 major phases. The participants were informed of the workshop goals and their task (Introduction), then they split into smaller groups to teach the dance clips to one another (Teaching), design a system to effectively teach the clip using what they learned from teaching one another (Designing) and then shared their systems with all the other groups (Presentation).
@@ 3. Method
    What are the current needs and design considerations of BLV learners seeking to learn dance asynchronously?
@@
    How can we design systems to make asynchronous dance instruction accessible through verbal narration, non-verbal audio, and/or haptics?
To incorporate the deep expertise and lived experience of people with disabilities and relevant experts, we selected to use co-design workshops — a participatory design approach that invites community members (e.g., BLV learners, dance instructors) to be creators and originators of the technology that they use rather than only end consumers (Kane et al., 2014; Magnusson et al., 2018). Similar to prior co-design workshops (Magnusson et al., 2018; Aflatoony and Lee, 2020; Valencia et al., 2021), we aimed to incorporate perspectives of not only potential end consumers of the technology (BLV dance learners) but also relevant experts in dance instruction (BLV and sighted dance instructors), and experts in multimodal technologies (audio description, sound design, haptics) to build on their complementary perspectives.
@@ 3.1. Workshop Design
We held three separate in-person co-design workshops that all featured the same five stages (Figure 2): (1) a pre-workshop survey to surface relevant individual dance backgrounds and lived experiences, (2) a large group introduction to share survey findings and discuss modalities for dance instruction, (3) small groups teaching example dances using non-visual techniques, then (4) designing systems for non-visual dance instruction using provided prototyping materials, and (5) presenting their systems to back to the large group for discussion. The four in-person workshop phases are shown in Figure 1. Each workshop was three hours long, and we compensated each participant \$90 for their participation. This study was approved by our institution’s Institutional Review Board (IRB).
@@
We began recruitment by reaching out to dance professors and students with teaching experience at our institution. We additionally connected with local BLV communities, and freelance sound producers and dance instructors. Furthermore, we partnered with an arts and accessibility organization and brought in two of their directors. Our three workshops included a total of 28 participants (16 female, 11 male, 1 non-binary), including 13 BLV participants (B1-B8b), 11 dance teachers (D1-D8b), two sound designers (S2, S6), a haptics engineer (H1), and an audio describer (A8) (Table 1). Many participants had combined expertise: 10 of 13 BLV participants had prior dance experience, B3b, D4a and H1 also had sound design expertise, and B2a, B3a, B7b, and B8a also had dance instruction expertise.
@@
Table 1. Participant demographic and background information, including gender, age, current vision, age of onset if blind, occupation, skill set (D = Dance, DI = Dance instruction, S = Sound design, H = Haptics engineering, A = Audio description), years of experience of the skill set, and dance styles if prior dance experience. Participant IDs are prefixed with “B” for BLV participants, or by primary expertise area if sighted. Numbers in participant IDs indicate the group number that the participant was a part of during Teaching and Designing stages.
@@ 3.1.2. Workshop Stages
We designed the workshop stages (Table 2) to maximize contributions by all workshop members, including using accessible workshop best practices (e.g., all materials accessible, ask for accommodations in advance, provide multiple opportunities to update access requests) and using a mix of individual, small group, and large group activities. We structured large group activities to encourage participation from every workshop member (e.g., think-pair-share, explicit turn order) and we selected small groups of 3-5 participants each (Table 2) to represent diverse expertise to foster complementary contributions.
Figure 2. Each individual workshop followed a five-stage process: 1) pre-workshop survey, 2) co-design workshop introduction, 3) teaching, 4) designing, and 5) presenting.
@@
From the survey results and our review of prior literature (Bläsing and Zimmermann, 2021; De Silva et al., 2023, 2025) we distilled three main themes for accessible asynchronous dance learning to address during the co-design workshop: (1) clarity, as visual demonstrations are often not fully supplemented by other sensory forms, (2) feedback, as remote learners are currently unable to have their movements corrected by visual comparison or touch, and (3) temporal and emotional qualities, as existing verbal descriptions insufficiently convey fluidity, weight, timing, emotions, and textures in movements that are important for learning dance.
@@
We first asked all participants and research staff to take turns introducing themselves to the group and answering the question “How would you describe what makes dance meaningful to you?” to consider aspects of dance they wanted to preserve in their later design exercise. We then shared survey themes around the benefits and challenges for accessible asynchronous movement education.
@@
Teaching & Designing: 85 minutes. Following the introductory presentation, we split participants into groups of 3-5 people with at least one BLV participant and one dance instructor per group (Table 2). We assigned each group a dance tutorial clip around 2.5 minutes long as a reference point for creating concrete designs of dance tutorial systems. We typically formed groups such that multiple participants shared experience in a common dance style, and we selected the tutorial clip to match the dance style experience of the group. The dance styles in the provided clips included Salsa, Contemporary, Hip-hop, Pop, and Cha-cha. As movement structures and teaching methods vary across different dance styles (Conroy and Bresnahan, 2025), the diversity of dance styles in our sample clips allowed for a wider range of teaching techniques across groups. We also ensured that selected clips were narrated and suitable for beginners. Partcipants remained with only their groups for the teaching and designing stages.
@@ 3 G8 B8a, B8b, D8a, D8b, A8 Cha-cha
Within their small groups, participants then discussed and designed a non-visual tutorial system for asynchronous dance education. We also asked participants to create a concrete prototype or a walkthrough of an accessible asynchronous dance instruction system using their given clip as an example. We provided participants with several prototyping materials to facilitate the designing stage (Figure 3): recording devices with raised buttons for recording verbal narration and non-verbal sound cues; a custom accessible soundboard application that allowed preview and adjustment of fluid/non-fluid music, natural sounds, and synthetic sounds (as informed by movement sonification literature (Frid et al., 2018; Landry and Jeon, 2020)); 3D movable human figures for demonstrating poses and haptic device placements; tactile stickers for prototyping haptic device placement on provided 3D figures or on participants themselves. Figure 4 shows participants using these prototyping tools to brainstorm and design their systems.
@@
Presenting: 30 minutes. Following group design sessions, all groups reconvened and were each asked to explain their non-visual tutorial system workflow in-depth, shown in Figure 5. We iterated on the workshop design following Workshop 1 and asked participants to additionally physically re-enact their system, using their given dance tutorial clip as an example. Each group designated one or more participants to act as the system, providing instruction, haptic and tactile feedback through (with permission) physical touch, and sound cues through the soundboard and onomatopoeia. Groups designated one or more participants to act as the learners from the system. Participants narrated system interactions during these re-enactments.
@@
3.1.3. Synchronous, In-Person Workshops for Asynchronous, Remote Dance Instruction Systems
Although our workshops addressed asynchronous and remote dance instruction, we held them in-person to allow for improved collaboration between participants through a shared physical space. The in-person setting allowed participants to establish common ground when learning and teaching one another example dances. Additionally, participants were able to use hands-on prototyping materials when designing their systems. Within their prototypes and during presentations, participants limited system interactions to what would be available in asynchronous contexts. For example, the use of touch was limited to where haptic devices would be placed, or verbal descriptions were limited to what would be provided by the system. Each group was accompanied by a member of the research team to center discussion on asynchronous contexts. Our in-person setup therefore allowed for engaging, collaborative workshops while addressing our problem space of asynchronous dance instruction.
@@ 4. Results
In this section, we share participants’ strategies and challenges of current dance instruction, their preferred approaches for using different modalities in remote instruction, and their concepts for future multimodal systems for end-to-end asynchronous dance instruction.
@@ 4.1. Current Strategies and Challenges
All groups watched the assigned example dance clips during the teaching phase to learn the sequence of movements, and all groups identified major challenges with directly following the videos alone: lack of context, lack of detail, and vague terminology. For example, G8’s clip directly began demonstrating the cha-cha routine without giving context such as how many steps and movements there were. All clips did not describe orientations of body parts with enough detail to recreate the movements. B8a expressed: “It’s not clear what step she is doing or if she is stepping at all. She could be stepping, or just shifting weight.”